An online music streaming service has a yearly membership fee of $50.88. There is an initial registration fee of $3.00 and the rest is paid monthly. To figure out how much each new member pays each month, solve for m in the equation shown, where m represents the monthly fee.

Answer by Theo(13342)

the yearly cost is 50.88
there is a 3 dollar registration fee and the rest is paid monthly.
50.88 minus 3 = 47.88
47.88 / 12 = 3.99

let y = the yearly cost.
let r = the registration fee.
let m = the monthly fee.
the equation is probably y = r + 12 * m where m represents the monthly fee and r represents the registration fee.

since r = 3 and m = 3.99, the equation becomes y = 3 + 12 * 3.99.
solve for y to get y = 50.88

if you started with the equation of y = r + 12 * m, and you knew that the yearly cost was 50.88 and you knew that the registration fee was 3 dollars, then the equation would have become:
50.88 = 3 + 12 * m
you would then have subtracted 3 from both sides of the equation to get:
50.88 - 3 = 12 * m
you would then have combined like terms to get:
47.88 = 12 * m
you would then have divided both sides of the equation by 12 to get:
47.88 / 12 = m
you would then have solved for m to get:
m = 47.88 / 12 = 3.99